Suppose that weekly demand for the electronic toy of a company is normally

distributed, with a mean of 2400 units and a standard deviation of 280 units.

i) What is the probability that the weekly demand for the electronic toy is

between 2350 and 2670?

ii) What is the value of a constant k, if


P(X > k) = 0.70


1
Expert's answer
2021-10-19T03:07:16-0400

i)


"P(2350<X<2670)=P(X<2670)-P(X\leq2350)"

"=P(Z<\dfrac{2670-2400}{280})-P(Z\leq\dfrac{2350-2400}{280})"

"\approx P(Z<0.9642857)-P(Z\leq-0.1785714)"

"\approx0.8325486-0.4291371\approx0.403412"


ii)


"P(X>k)=1-P(X\leq k)"

"=1-P(Z\leq\dfrac{k-2400}{280})=0.7"

"P(Z\leq\dfrac{k-2400}{280})=0.3"

"\dfrac{k-2400}{280}\approx-0.5244005"

"k\approx2253"